The equation of a line is y= -1/2x -3
Complete the table for this equation.
x y (x,y)
? -1 (?,?)
-2 ? (?,?)
? -4 (?,?)
6 ? (?,?)
8 ? (?,?)

Answer by rfer(12655)

-1=-1/2x-3
+2=-1/2x
-4=x
(-4,-1)
------------
you try the others